OpenStudy (anonymous):

Given that y varies directly with x, write the linear function relating the two variables if y= 4 when x =36

OpenStudy (anonymous):

y is proportional to x thus u can express it as y=kx where k is a constant Substitute y=4 and x= 36 to find out what k is 4 = k(36) k = 1/9 Linear function: y= x/9